OpenStudy (mm3591):

Use the substitution method to solve the system of equations. Enter your answer as an ordered pair. y = 2x + 1 y = 6x - 1

jimthompson5910 (jim_thompson5910):

y = 2x + 1 6x-1 = 2x + 1 ... replace y with 6x-1 now solve for x and tell me what you get

OpenStudy (mm3591):

1/3

jimthompson5910 (jim_thompson5910):

incorrect

jimthompson5910 (jim_thompson5910):

6x-1 = 2x + 1 6x-1+1 = 2x + 1+1 ... Add 1 to both sides. 6x = 2x + 2 6x - 2x = 2x + 2 - 2x ... Subtract 2x from both sides. 4x = 2 x = ???

OpenStudy (mm3591):

2

jimthompson5910 (jim_thompson5910):

x = 2 does NOT make 4x = 2 true

jimthompson5910 (jim_thompson5910):

4x = 2 4*2 = 2 ... replace x with 2 8 = 2 ... this is false

OpenStudy (mm3591):

Thanks

jimthompson5910 (jim_thompson5910):

if 4x = 2, then what must x be?

OpenStudy (mm3591):

I don't even know anymore

jimthompson5910 (jim_thompson5910):

divide both sides by 4

jimthompson5910 (jim_thompson5910):

2/4 = ???

OpenStudy (mm3591):

0.5

jimthompson5910 (jim_thompson5910):

yep or 1/2

jimthompson5910 (jim_thompson5910):

now that you know x = 0.5, you can use this to find y

jimthompson5910 (jim_thompson5910):

by plugging that x value into any equation with x & y in it, then solving for y

OpenStudy (mm3591):

Thanks

jimthompson5910 (jim_thompson5910):

you're welcome